# A QUBO Formulation for the Generalized LinkedIn Queens and Takuzu/Tango Game

### v2

@@ 2.6 Chess Piece Problems
    $\sum_{(i,j)\in R^{\prime}_{k}}x_{ij}=1\quad\forall k\in[0,N^{\prime}_{R}-1]$
@@
$Q_{g}=\sum_{k=0}^{N^{\prime}_{R}-1}\left(1-\sum_{(i,j)\in R^{\prime}_{k}}x_{ij}\right)^{2}$ (23)
@@
$Q=$ $Q_{g}+Q_{t}=\sum_{k=0}^{N^{\prime}_{R}-1}\left(1-\sum_{(i,j)\in R^{\prime}_{k}}x_{ij}\right)^{2}+$
$+\sum_{i,j=0}^{N-1,M-1}\sum_{(l,m)\in C\left(M_{P_{ij}}\right)_{ij}}x_{ij}x_{lm}.$ (25)
@@
Given a board of arbitrary shape, of maximum height $N$ and maximum width $M$, on which there are placed a number $N_{p}$ of initial pieces in positions given by the set $A_{I}$, the objective will be to place the maximum number possible of pieces on the board such that no piece threatens another. Each cell $(i,j)$ can be empty or have the piece of type $P_{ij}$, with possible moves $M_{P_{ij}}$, and can reach cells $C\left(M_{P_{ij}})_{ij}\right)$. A solution of this problem is represented in Fig. 5 b.
Since we will only keep the threat constraint, we can eliminate the coloured term, and we have to include a new term that reduces the energy the more tiles there are on the board. This counting term is
$Q_{n}=-\sum_{i,j=1}^{N-1,M-1}x_{ij}$ (26)
so that the energy will decrease by one unit for each piece added to the board. This formulation even allows a different weight to be associated with each type of piece, so that it is more important to place some types of pieces than others.
To get the total QUBO, if we add directly the terms $Q_{t}$ and $Q_{m}$ we can have cases in which, by adding a single piece that threatens only one piece, we will have the same energy. Therefore, to make sure that the constraint is met, we will multiply the $Q_{t}$ term by a constant $\lambda$ large enough so that it is not convenient to add pieces that do not meet the constraint. Therefore, the total QUBO is
$Q=$ $Q_{n}+\lambda Q_{t}=-\sum_{i,j=1}^{N-1,M-1}x_{ij}+$
$+$ $\lambda\sum_{i,j=0}^{N-1,M-1}\sum_{(l,m)\in C\left(M_{P_{ij}}\right)_{ij}}x_{ij}x_{lm}.$ (27)
@@ 3.1 Problem definition
In this case, we can see that if we replace the suns by zeros and the moons by ones, and allow the board to be $N\times M$, we will have a generalization of Takuzu. We call this problem the Takuzu/Tango Problem (TTP).
To formulate this problem, the state of a board is given by a set of variables $x_{ij}$, which indicate the value in the cell of row $i$ and column $j$. Therefore, the constraints of the problem can be expressed as:
@@
  • Vertical repetition condition: there cannot be more than two zeros or ones in a row vertically. Therefore, for each group of 3 vertically consecutive cells, there can only be 1 or 2 ones, which implies that their associated variables can only add up to 1 or 2 (Fig. 7 a):
@@
  • Horizontal repetition condition: there cannot be more than two zeros or ones in a row horizontally. Therefore, for each group of 3 horizontally consecutive cells, there can only be 1 or 2 ones, which implies that their associated variables can only add up to 1 or 2 (Fig. 7 b):
    $\sum_{k=j}^{j+2}$ $x_{ik}\in\{1,2\}$ (31)
    $\forall i$ $\in[1,N],j\in[1,M-2].$
  • Equal sign condition: If there is an ’=’ symbol between two cells, both cells must have the same value. These symbols can be between two cells horizontally or between two cells vertically (Fig. 7 d). There are $N_{He}$ horizontal symbols and $N_{Ve}$ vertical symbols. The $k$-th horizontal ’=’ is between positions $(I^{He}_{k},J^{He}_{k})$ and $(I^{He}_{k},J^{He}_{k}+1)$, while the $k$-th vertical ’=’ is between positions $(I^{Ve}_{k},J^{Ve}_{k})$ and $(I^{Ve}_{k}+1,J^{Ve}_{k})$. Therefore, for the horizontals, the following is satisfied
    $x_{I^{He}_{k},J^{He}_{k}}=x_{I^{He}_{k},J^{He}_{k}+1}\ \forall k\in[1,N_{He}],$ (32)
    and for the verticals
    $x_{I^{Ve}_{k},J^{Ve}_{k}}=x_{I^{Ve}_{k}+1,J^{Ve}_{k}}\quad\forall k\in[1,N_{Ve}].$ (33)
  • Cross sign condition: If there is an ’x’ symbol between two cells, both cells must have different value. These symbols can be between two cells horizontally or between two cells vertically (Fig. 7 e). There are $N_{Hc}$ horizontal symbols and $N_{Vc}$ vertical symbols. The $k$-th horizontal ’x’ is between positions $(I^{Hc}_{k},J^{Hc}_{k})$ and $(I^{Hc}_{k},J^{Hc}_{k}+1)$, while the $k$-th vertical ’x’ is between positions $(I^{Vc}_{k},J^{Vc}_{k})$ and $(I^{Vc}_{k}+1,J^{Vc}_{k})$. Therefore, for the horizontals, the following is satisfied
    $x_{I^{Hc}_{k},J^{Hc}_{k}}\neq x_{I^{Hc}_{k},J^{Hc}_{k}+1}\ \forall k\in[1,N_{Hc}],$ (34)
    and for the verticals
    $x_{I^{Vc}_{k},J^{Vc}_{k}}\neq x_{I^{Vc}_{k}+1,J^{Vc}_{k}}\quad\forall k\in[1,N_{Vc}].$ (35)
  • Global non-repetition constraint: no two rows can have the same combination of zeros and ones. As the positions of the ones condition those of the zeros, we will only have to consider those of the ones. If two cells in the same row are at one, the product of their values will be one. In any other case it will be zero. Therefore, all the elements of two rows will coincide if the sum of their products is equal to the number of ones in each row.
    $\sum_{j=1}^{M}x_{aj}x_{bj}<\frac{M}{2}\ \forall a,b\in[1,N]\ |\ a\neq b.$ (36)
    Neither can there be two columns with the same combination of zeros and ones.
    $\sum_{i=1}^{N}x_{ia}x_{ib}<\frac{N}{2}\ \forall a,b\in[1,M]\ |\ a\neq b.$ (37)
@@ 3.2 QUBO Formulation
The QUBO formulation of this problem is composed of different terms that are responsible for imposing each of the constraints. Before starting to formulate each term, we will discuss how to optimize the number of variables to be determined by fixing trivial values for some variables given an initialization. This way, some variables will have a fixed value in the QUBO terms, but they will not be introduced as variables to be optimized in algorithms such as QAOA or Quantum Annealers, but will be part of the coefficients of the cost matrix. This make it possible to reduce the resources, e.g. qubits, to be used in quantum devices.
If we have a set of zeros and ones placed in their positions $A_{I0}$ and $A_{I1}$, the variables associated with those cells do not need to be determined. This will reduce the number of variables to optimize to $NM-N_{I0}-N_{I1}$.
In case any of these already fixed variables is adjacent to an ’=’ symbol, we will trivially fix the adjacent variable to its same value. If the adjacent symbol is an ’x’, we will fix the adjacent variable to the opposite value. This allow us to remove those variables from the optimization in the same way.
In case a row or column has already fixed as many zeros or ones as it should have by the regularity condition, the rest of the variables are set to the opposite value, as this is the only way to satisfy this constraint. With this we can eliminate even more variables trivially.
In case there are 2 adjacent variables fixed with the same value, we will automatically know that the ones adjacent to both of them that follow in their line have to have the opposite value, which allows us to fix some more variables.
Once the number of variables to be determined has been optimized, we proceed with the creation of the QUBO terms. The first term we need to create is the horizontal repetition term, which is
$Q_{HR}=\sum_{i,j=1}^{N,M-2}\left(\frac{3}{2}-\sum_{k=j}^{j+2}x_{ik}\right)^{2}$ (38)
where we have that the minimum is centered between 1 and 2 using a fractional QUBO term [13], so that in each horizontal group of 3 variables the minimum value of the energy ($\frac{1}{4}$) is given when 1 or 2 variables have the value 1. For the vertical condition, we have an analogous term
@@
For the regularization conditions, we only need a term that imposes that the sum of the variables per row or column results in a specific value. These terms are for each case
@@
Finally, we will have the 4 terms necessary to impose the constraints of ’=’ horizontal, ’=’ vertical, ’x’ horizontal and ’x’ vertical. For the ’=’ terms, we need the energy when they are equal to be zero. We achieve this by knowing that, for a binary case $x(1-y)+(1-x)y$ is null if $x=y$ and 1 if $x\neq y$. We need the second summand so that there is no null if $x=0$ and $y=1$. Thus, the terms for equals are
$Q_{He}=\sum_{k=1}^{N_{He}}$ $\left(x_{I_{k}^{He},J_{k}^{He}}\left(1-x_{I_{k}^{He},J_{k}^{He}+1}\right)+\right.$
$+$ $\left.\left(1-x_{I_{k}^{He},J_{k}^{He}}\right)x_{I_{k}^{He},J_{k}^{He}+1}\right),$ (42)
$Q_{Ve}=\sum_{k=1}^{N_{Ve}}$ $\left(x_{I_{k}^{Ve},J_{k}^{Ve}}\left(1-x_{I_{k}^{Ve}+1,J_{k}^{Ve}}\right)+\right.$
$+$ $\left.\left(1-x_{I_{k}^{Ve},J_{k}^{Ve}}\right)x_{I_{k}^{Ve}+1,J_{k}^{Ve}}\right).$ (43)
For the case ’x’ we need a term that is null only when both values are different. To do so, we will take advantage of the fact that $xy+(1-x)(1-y)$ is equal to 0 if $x\neq y$ and 1 if $x=y$. Thus, the terms are
$Q_{Hc}=\sum_{k=1}^{N_{Hc}}$ $\left(x_{I_{k}^{Hc},J_{k}^{Hc}}x_{I_{k}^{Hc},J_{k}^{Hc}+1}+\right.$
$+$ $\left.\left(1-x_{I_{k}^{Hc},J_{k}^{Hc}}\right)\left(1-x_{I_{k}^{Hc},J_{k}^{Hc}+1}\right)\right),$ (44)
$Q_{Vc}=\sum_{k=1}^{N_{Vc}}$ $\left(x_{I_{k}^{Vc},J_{k}^{Vc}}x_{I_{k}^{Vc}+1,J_{k}^{Vc}}+\right.$
$+$ $\left.\left(1-x_{I_{k}^{Vc},J_{k}^{Vc}}\right)\left(1-x_{I_{k}^{Vc}+1,J_{k}^{Vc}}\right)\right).$ (45)
Given this formulation, we can perform the following optimization of variables by taking advantage of the symbols ’=’ and ’x’. As we know, if there is one of these symbols between two cells, both variables must have the same or different value. In this case in which we can only have two possible values, this implies that these two variables are going to be completely dependent, so that if we determine one, the other is automatically determined. Therefore, these last four terms of the formulation are unnecessary if we reduce the number of variables as follows.
In our optimization, we optimize only one of the variables of the pair affected by the symbols, and in case we have a chain of variables all affected by each other, we optimize only one of them. The rule we follow is that in each horizontal pair, we optimize the variable on the left, and in each vertical pair, the one on top. In case of having whole regions affected, we choose the variable that is more to the left, and in case of having several in the same column, the one that is higher among them.
To take this dependence into account, we will make the following substitution in the first four terms:
@@
In this way, we are able to impose this constraint automatically, without having to optimize it, reducing the number of variables to optimize. Thus, the number of variables to be optimized will be, at the most
$NM-N_{I0}-N_{I1}-N_{He}-N_{Hc}-N_{Ve}-N_{Vc}.$
Finally, the global non-repetition condition cannot be written as a QUBO term, so we will have to be confident that the solutions to the problem do not conflict with this constraint.
Therefore, the QUBO formulation of the TTP is
$Q=$ $Q_{HR}+Q_{VR}+Q_{Hr}+Q_{Vr}=$
$=$ $\sum_{i,j=1}^{N,M-2}\left(\frac{3}{2}-\sum_{k=j}^{j+2}x_{ik}\right)^{2}+$
$+$ $\sum_{i,j=1}^{N-2,M}\left(\frac{3}{2}-\sum_{k=i}^{i+2}x_{kj}\right)^{2}+$
$+$ $\sum_{i=1}^{N}\left(\frac{M}{2}-\sum_{j=1}^{M}x_{ij}\right)^{2}+$
$+$ $\sum_{j=1}^{M}\left(\frac{N}{2}-\sum_{i=1}^{N}x_{ij}\right)^{2}.$ (50)
Thus, a solution to the problem that satisfies the constraints will have an energy of $\frac{1}{2}(NM-N-M)$.

### v3

@@ 2.6 Chess Piece Problems
    $\sum_{(i,j)\in R^{\prime}_{k}}x_{ij}=1\quad\forall k\in[1,N^{\prime}_{R}]$
@@
$Q_{g}=\sum_{k=1}^{N^{\prime}_{R}}\left(1-\sum_{(i,j)\in R^{\prime}_{k}}x_{ij}\right)^{2}$ (36)
@@
$Q=Q_{g}+Q_{t}=\sum_{k=1}^{N^{\prime}_{R}}\left(1-\sum_{(i,j)\in R^{\prime}_{k}}x_{ij}\right)^{2}+\sum_{i=1}^{N}\sum_{j=1}^{M}\sum_{(l,m)\in C\left(M_{P_{ij}}\right)_{ij}}x_{ij}x_{lm}.$ (38)
@@
Given a board of arbitrary shape, of maximum height $N$ and maximum width $M$, on which there are placed a number $N_{p}$ of initial pieces in positions given by the set $A_{I}$, the objective will be to place the maximum number possible of pieces on the board such that no piece threatens another. Each cell $(i,j)$ can be empty or have the piece of type $P_{ij}$, with possible moves $M_{P_{ij}}$, and can reach cells $C\left(M_{P_{ij}}\right)_{ij}$. A solution of this problem is represented in Fig. 5 b.
Since we only keep the threat constraint, we eliminate the coloured term and add a reward term that favors occupied cells. In weighted form, let $w_{ij}\geq 0$ be the reward for placing a piece in cell $(i,j)$, and define
$Q_{n}=-\sum_{i=1}^{N}\sum_{j=1}^{M}w_{ij}x_{ij}.$ (39)
The unweighted case corresponds to $w_{ij}=1$ for all cells.
Therefore, the total QUBO is
$Q=Q_{n}+\lambda Q_{t}=-\sum_{i=1}^{N}\sum_{j=1}^{M}w_{ij}x_{ij}+\lambda\sum_{i=1}^{N}\sum_{j=1}^{M}\sum_{(l,m)\in C\left(M_{P_{ij}}\right)_{ij}}x_{ij}x_{lm}.$ (40)
Proposition 2.3.
If $\lambda>\max_{i,j}w_{ij}$, then no global minimizer of the Max Chess Pieces QUBO contains a conflicting pair of pieces. In particular, in the unweighted case it is sufficient to take $\lambda>1$.
Proof.
Assume a minimizer contains a conflicting pair. Then there exists an occupied cell $z=(i,j)$ belonging to at least one conflict edge. If we switch $x_{z}$ from $1$ to 0, the reward term increases by at most $w_{ij}\leq\max_{i,j}w_{ij}$, while the conflict term decreases by at least $\lambda$ because at least one conflicting product disappears. Since $\lambda>\max_{i,j}w_{ij}$, the total energy decreases, contradicting optimality. ∎
@@ 3.1 Problem definition
In this case, if we replace the suns by zeros and the moons by ones, and allow the board to be $N\times M$, we obtain a common local framework for both Takuzu and Tango. We call this local framework the Takuzu/Tango Problem (TTP): it contains the clues, the row and column balance conditions, the prohibition of three equal consecutive values, and optionally the ’=’ and ’x’ relations. Classical Takuzu adds one more global condition, namely that no two rows and no two columns can coincide, whereas Tango does not require this global uniqueness.
To formulate the local TTP, the state of a board is given by variables $x_{ij}\in\{0,1\}$, which indicate the value in the cell of row $i$ and column $j$. Therefore, the local constraints can be expressed as:
@@
  • Vertical repetition condition: there cannot be more than two zeros or ones in a row vertically, so every 3-cell vertical window must contain one or two ones (Fig. 7 a):
@@
  • Horizontal repetition condition: there cannot be more than two zeros or ones in a row horizontally, so every 3-cell horizontal window must contain one or two ones (Fig. 7 b):
    $\sum_{k=j}^{j+2}x_{ik}\in\{1,2\}\ \forall i\in[1,N],\ j\in[1,M-2].$ (44)
  • Equal sign condition: if there is an ’=’ symbol between two cells, both cells must have the same value.
  • Cross sign condition: if there is an ’x’ symbol between two cells, both cells must have different values.
For classical Takuzu, one must additionally impose the global non-repetition condition
$(x_{r1},\ldots,x_{rM})\neq(x_{s1},\ldots,x_{sM})\quad\forall 1\leq r<s\leq N,$ (45)
and analogously
$(x_{1c},\ldots,x_{Nc})\neq(x_{1t},\ldots,x_{Nt})\quad\forall 1\leq c<t\leq M.$ (46)
These global uniqueness constraints will be treated separately below: the local QUBO is exact for Tango and for the Takuzu/Tango variant without global uniqueness, while an exact QUBO for full Takuzu requires additional auxiliary variables.
@@ 3.2 QUBO Formulation
The local QUBO formulation is composed of terms that impose the clues, the balance conditions, the prohibition of three consecutive equal values and, when present, the parity signs. We first discuss how to reduce trivial variables. If we have zeros and ones already placed in positions $A_{I0}$ and $A_{I1}$, those cells do not need to be optimized. Moreover, if one of these fixed values is linked by an ’=’ or ’x’ sign to another cell, then the adjacent cell can also be fixed immediately. The same happens when a row or column already contains as many zeros or ones as allowed by the balance condition, or when two equal fixed adjacent values force the third value in a length-3 window.
In unreduced form, the clue penalty can be written as
$Q_{\mathrm{clues}}=\sum_{(i,j)\in A_{I0}}x_{ij}+\sum_{(i,j)\in A_{I1}}(1-x_{ij}),$ (47)
which is a QUBO because the terms are linear in binary variables. After preprocessing, this term becomes a constant and may be omitted from the reduced optimization problem.
The horizontal repetition term is
$Q_{HR}=\sum_{i,j=1}^{N,M-2}\left(\frac{3}{2}-\sum_{k=j}^{j+2}x_{ik}\right)^{2},$ (48)
and the vertical analogue is
@@
These are fractional QUBO terms of the form discussed in (3): their minimizers are exactly the assignments for which each 3-cell window contains one or two ones, but each satisfied window contributes $1/4$ rather than 0. If zero energy on each satisfied window is desired, one may subtract that constant or replace the term by the equivalent zero-based penalty from the preliminaries.
For the regularity conditions we use
@@
For the sign constraints we can directly use the equality and difference penalties from the preliminaries:
$Q_{He}=\sum_{k=1}^{N_{He}}\left(x_{I_{k}^{He},J_{k}^{He}}-x_{I_{k}^{He},J_{k}^{He}+1}\right)^{2},$ (52)
$Q_{Ve}=\sum_{k=1}^{N_{Ve}}\left(x_{I_{k}^{Ve},J_{k}^{Ve}}-x_{I_{k}^{Ve}+1,J_{k}^{Ve}}\right)^{2},$ (53)
$Q_{Hc}=\sum_{k=1}^{N_{Hc}}\left(x_{I_{k}^{Hc},J_{k}^{Hc}}+x_{I_{k}^{Hc},J_{k}^{Hc}+1}-1\right)^{2},$ (54)
$Q_{Vc}=\sum_{k=1}^{N_{Vc}}\left(x_{I_{k}^{Vc},J_{k}^{Vc}}+x_{I_{k}^{Vc}+1,J_{k}^{Vc}}-1\right)^{2}.$ (55)
The ’=’ and ’x’ symbols are naturally described by a parity graph $G=(V,E)$ whose vertices are cells and whose edges carry a parity $p_{e}\in\{0,1\}$: $p_{e}=0$ for ’=’ and $p_{e}=1$ for ’x’. The constraint on an edge $e=(u,v)$ is
$x_{u}\oplus x_{v}=p_{e}.$ (56)
Proposition 3.1.
The parity system on $G$ is consistent if and only if the XOR of the parities along every cycle of $G$ is 0.
Proof.
If one XORs the equations $x_{u}\oplus x_{v}=p_{e}$ along a cycle, every vertex variable appears twice and cancels, so any satisfying assignment must yield XOR 0 for the edge parities. Conversely, if every cycle has XOR 0, fixing one representative value in each connected component determines all other vertices uniquely by path parity, and the cycle condition guarantees that this value does not depend on the chosen path. ∎
Therefore the sign constraints can be reduced by a union-find structure with parity. If the system is inconsistent, the puzzle has no solution. If it is consistent, every connected component $C$ of the parity graph admits a representative variable $y_{C}$ and fixed parities $\pi_{v}\in\{0,1\}$ such that
$x_{v}=y_{C}\oplus\pi_{v}\quad\forall v\in C.$ (57)
If a clue fixes one vertex in the component, then it also fixes $y_{C}$; if two clues induce incompatible values, the puzzle has no solution. Consequently, the exact number of free variables after reducing the signs is the number of parity-connected components that are not fixed by clues. The expression
$NM-N_{I0}-N_{I1}-N_{He}-N_{Hc}-N_{Ve}-N_{Vc}$
is only a crude upper bound when every sign contributes an independent reduction; in general it is not exact because cycles and redundant constraints do not reduce the number of free variables further.
At the level of a single edge, the previous reduction can still be written through the substitutions
@@
These local substitutions are single-edge instances of the parity-component reduction described above.
Hence the local TTP QUBO can be written as
$Q_{TTP}^{\mathrm{local}}=\ Q_{\mathrm{clues}}+Q_{HR}+Q_{VR}+Q_{Hr}+Q_{Vr}+Q_{He}+Q_{Ve}+Q_{Hc}+Q_{Vc}.$ (62)
For Tango, or for the Takuzu/Tango variant without global row/column uniqueness, this formulation is exact after a consistent parity preprocessing. For classical Takuzu it is only a relaxation, because it does not yet enforce that the completed rows and columns be pairwise distinct. If one keeps the fractional forms $Q_{HR}$ and $Q_{VR}$, every feasible local board has the constant offset (NMNM)/2(NM-N-M)/2 coming from the length-3 windows; subtracting this constant does not change the minimizers.
3.2.1 Exact Takuzu QUBO with witness variables
We now restrict to the classical Takuzu board, which is square of size $N\times N$, and add an exact QUBO for the global non-repetition constraint without using slack variables.
For each pair of rows $r<s$ and each column $j$, introduce witness variables
$u^{01}_{rsj},u^{10}_{rsj}\in\{0,1\}.$
The intended meaning is that $u^{01}_{rsj}=1$ certifies $x_{rj}=0$ and $x_{sj}=1$, while $u^{10}_{rsj}=1$ certifies $x_{rj}=1$ and $x_{sj}=0$. Define
$S^{\mathrm{row}}_{rs}=\sum_{j=1}^{N}\left(u^{01}_{rsj}+u^{10}_{rsj}\right).$ (63)
Then we add the penalty
$Q^{\mathrm{row}}_{\neq}=A_{\neq}\sum_{1\leq r<s\leq N}\Bigg[\left(S^{\mathrm{row}}_{rs}-1\right)^{2}+\sum_{j=1}^{N}\Big(u^{01}_{rsj}x_{rj}+u^{01}_{rsj}(1-x_{sj})+u^{10}_{rsj}(1-x_{rj})+u^{10}_{rsj}x_{sj}\Big)\Bigg].$ (64)
The term $\left(S^{\mathrm{row}}_{rs}-1\right)^{2}$ forces exactly one active witness for each pair of rows, while the consistency terms force that an active witness must point to a column in which the two rows genuinely differ.
Proposition 3.2.
For fixed board variables $x$, there exists an assignment of the row-witness variables $u$ with $Q^{\mathrm{row}}_{\neq}=0$ if and only if all rows are pairwise distinct.
Proof.
Assume rows $r$ and $s$ are different. Then there exists a column $j$ with $x_{rj}\neq x_{sj}$. If $(x_{rj},x_{sj})=(0,1)$, set $u^{01}_{rsj}=1$; if $(x_{rj},x_{sj})=(1,0)$, set $u^{10}_{rsj}=1$; keep all other witnesses for that pair at 0. Then $S^{\mathrm{row}}_{rs}=1$ and every consistency term vanishes. Conversely, if $Q^{\mathrm{row}}_{\neq}=0$, then for each pair $r<s$ exactly one witness is active. If it is $u^{01}_{rsj}$, the consistency terms force $x_{rj}=0$ and $x_{sj}=1$; if it is $u^{10}_{rsj}$, they force $x_{rj}=1$ and $x_{sj}=0$. In both cases the rows differ. ∎
Similarly, for each pair of columns $c<t$ and each row $i$, introduce
$v^{01}_{cti},v^{10}_{cti}\in\{0,1\},$
where $v^{01}_{cti}=1$ certifies $x_{ic}=0$ and $x_{it}=1$, and $v^{10}_{cti}=1$ certifies $x_{ic}=1$ and $x_{it}=0$. Define
$S^{\mathrm{col}}_{ct}=\sum_{i=1}^{N}\left(v^{01}_{cti}+v^{10}_{cti}\right),$ (65)
and add
$Q^{\mathrm{col}}_{\neq}=A_{\neq}\sum_{1\leq c<t\leq N}\Bigg[\left(S^{\mathrm{col}}_{ct}-1\right)^{2}+\sum_{i=1}^{N}\Big(v^{01}_{cti}x_{ic}+v^{01}_{cti}(1-x_{it})+v^{10}_{cti}(1-x_{ic})+v^{10}_{cti}x_{it}\Big)\Bigg].$ (66)
Proposition 3.3.
For fixed board variables $x$, there exists an assignment of the column-witness variables $v$ with $Q^{\mathrm{col}}_{\neq}=0$ if and only if all columns are pairwise distinct.
Proof.
If columns $c$ and $t$ are different, choose one row $i$ in which they differ and activate the corresponding witness $v^{01}_{cti}$ or $v^{10}_{cti}$. Then $S^{\mathrm{col}}_{ct}=1$ and all consistency terms vanish. Conversely, if $Q^{\mathrm{col}}_{\neq}=0$, each pair $c<t$ has exactly one active witness, and that witness certifies a row in which the two columns differ. ∎
Let
$Q_{\mathrm{balance}}=Q_{Hr}+Q_{Vr},$ (67)
$Q_{\mathrm{triple}}=Q_{HR}+Q_{VR}.$ (68)
Then an exact QUBO for classical Takuzu is
$Q_{Takuzu}^{\mathrm{exact}}=Q_{\mathrm{clues}}+Q_{\mathrm{balance}}+Q_{\mathrm{triple}}+Q^{\mathrm{row}}_{\neq}+Q^{\mathrm{col}}_{\neq}.$ (69)
If one keeps the fractional form in $Q_{\mathrm{triple}}$, the feasible boards share a constant energy offset; the minimizing boards are nevertheless unchanged, and one may subtract the constant or use the zero-based triple penalty from the preliminaries if a zero feasible energy is preferred.
Theorem 3.1.
Minimizing $Q_{Takuzu}^{\mathrm{exact}}$ over the board variables and the witness variables produces exactly the solutions of the full Takuzu problem.
Proof.
The term $Q_{\mathrm{clues}}$ enforces the clues, $Q_{\mathrm{balance}}$ enforces the row and column balance conditions, and $Q_{\mathrm{triple}}$ enforces that no horizontal or vertical length-3 window is all-zero or all-one. By the two previous propositions, $Q^{\mathrm{row}}_{\neq}$ enforces that no two rows coincide and $Q^{\mathrm{col}}_{\neq}$ enforces that no two columns coincide. Therefore every minimizer satisfies exactly the Takuzu rules. Conversely, if a board is a valid Takuzu solution, then every pair of rows and every pair of columns differ somewhere, so one can activate one consistent witness for each pair and obtain zero contribution from the new uniqueness terms; hence valid Takuzu boards are precisely the minimizing configurations. ∎
The row witnesses use $2N\binom{N}{2}=N^{2}(N-1)$ auxiliary variables, and the column witnesses use the same amount. Therefore the exact Takuzu construction requires
$2N^{2}(N-1)=O(N^{3})$
auxiliary binary variables in total, and no slack variables are needed for the row/column uniqueness constraint. This shows that the global uniqueness condition can indeed be incorporated into an exact QUBO formulation, at the cost of auxiliary witness variables.